OpenStudy (anonymous):

Small question.. 49/42 simplified? Also wondering if I got this right: 7/7 times 7/6= 49/42 (without simplifying)

OpenStudy (mathstudent55):

Divide both the numerator and denominator by the GCF of 49 and 42. I think you already know it's 7. You are right.

OpenStudy (anonymous):

Thank you!

OpenStudy (broskishelleh):

Yes they both equal 1.1666666... But that they are equal you are right

OpenStudy (mathstudent55):

\(\dfrac{49}{42} = \dfrac{7 \times 7}{7 \times 6} = \dfrac{7}{7} \times \dfrac{7}{6} = 1 \times \dfrac{7}{6} = \dfrac{7}{6}\)

OpenStudy (broskishelleh):

Thank you @mathstudent55 I am ashamed of getting that wrong :(